Ms. Ratchford wants to go for a walk. She travels to Starbucks which is 3.5 km [N11W] she then goes to Walmart to get groceries and travels 1.5 KM[S45E]. What is her displacement?


1
Expert's answer
2021-06-25T10:45:15-0400

"\text{let:}"

"\text{point O starting point of movement}"

"\text{point S - Starbucks location}"

"\text{point W - Walmart location}"

"\alpha=\angle(\vec {OS},\vec {SW})"

"\alpha= [N11W] +\angle({\vec{N}},{\vec{E}})+ [S45E]=146\degree"

"\vec {OW}= \vec {OS}+\vec {SW}"

"|\vec {OW}|=\sqrt{|\vec {OS}|^2+|\vec {SW}|^2+2*|\vec {OS}|*|\vec {SW}|*\cos\alpha}"

"|\vec {OW}|=\sqrt{3.5^2+1.5^2+2*1.5*3.5*\cos146\degree}=2.4"

"\text{Aswer: displacement = 2.4 km}"